In a local bar, a customer slides an empty beer mug down the counter for a refill. The height of the counter is 1.46 m. The mug slides off the counter and strikes the floor 2.00 m from the base of the counter.
What was the direction of the mug's velocity just before it hit the floor?
(Below the horizontal)
Answer #1

initial velocity along y axis=0

so y= Voy t+ 1/2 gt2

1.46=0(t)+1/2 (9.8)t2=0.55 seconds

let Vx= initial velocity along horzontal

Vx= x/t

Vx=2/(0.55)=3.66 m/s

now Vy= Voy+ gt

Vy=0+ 9.8(0.55)=5.39 m/s

theta= tan-1(Vy/Vx)

theta=tan-1(5.39/3.66)

theta=55.8o